# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v1

@@ Abstract
The predominant approach to facial action unit (AU) detection revolves around a supervised multi-label binary classification problem. Existing methodologies often encode pixel-level information of AUs, thereby imposing substantial demands on model complexity and expressiveness. Moreover, this practice elevates the susceptibility to overfitting due to the presence of noisy AU labels. In the present study, we introduce a contrastive learning framework enhanced by both supervised and self-supervised signals. The objective is to acquire discriminative features, deviating from the conventional pixel-level learning paradigm within the domain of AU detection. To address the challenge posed by noisy AU labels, we augment the supervised signal through the introduction of a self-supervised signal. This augmentation is achieved through positive sample sampling, encompassing three distinct types of positive sample pairs. Furthermore, to mitigate the imbalanced distribution of each AU type, we employ an importance re-weighting strategy tailored for minority AUs. The resulting loss, denoted as AUNCE, is proposed to encapsulate this strategy. Our experimental assessments, conducted on two widely-utilized benchmark datasets (BP4D and DISFA), underscore the superior performance of our approach compared to state-of-the-art methods in the realm of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
@@ I Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding. It has introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2], enabling their simulation. In contemporary times, AU detection [3, 4, 5, 6] has evolved into a technique utilizing computer vision algorithms to automatically identify and track AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction [7], emotion analysis [8, 9], and monitoring in activities such as car driving [10].
Refer to caption (a)
Refer to caption (b)
Refer to caption (c)
Refer to caption (d)
Figure 1: Illustration of our proposed positive sample sampling strategy. From left to right: (a) original images, (b) self-supervised images obtained by simple augmentations, (c) images with the highest similarity to (a), and (d) images with the lowest similarity to (a). We use (a) with (b), (c), mixture of all positive samples as our positive sample pairs.
In recent years, a predominant trend in supervised frameworks for AU detection has been observed, as evidenced by various studies [11, 12, 13, 14, 15]. These frameworks have predominantly employed the traditional binary cross-entropy loss, which serves as a metric for gauging the disparity between predicted and actual ground truth labels, providing a systematic means to guide the learning process. However, encoding the pixel-level information is notable for its heightened sensitivity to models characterized by a substantial number of parameters, as it endeavors to encode information at the pixel level within the input space. This imposes a notable demand on the fitting capability of the models. Additionally, pixel-level generative learning, while computationally intensive, may prove superfluous for AU detection tasks, given their nature as multi-label binary classification problems. The pursuit of pixel-level features not only introduces unwarranted model complexity but also augments the susceptibility to overfitting.
@@
Representation learning approaches rooted in the discriminative contrastive learning paradigm have demonstrated substantial promise, yielding state-of-the-art outcomes [16, 17, 18, 19, 20]. The fundamental concept of contrastive learning involves maximizing the consistency of positive example features that are semantically similar while minimizing the consistency of negative example features that are semantically dissimilar [21]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized, relieving the model from encoding pixel-level information [22]. However, the direct integration of contrastive learning into AU detection tasks encounters two principal challenges. The first challenge pertains to the noisy label problem [23]. Due to the weak amplitude of AU occurrence and the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit errors and noisy labels during the labeling process. Consequently, fitting models with noisy and erroneous samples significantly undermines the overall fitting ability of models. The second challenge is the class imbalance problems among different AUs [11]. Most AU classes manifest a substantial imbalance between positive and negative samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance not only diminishes the fitting efficiency of neural networks but also influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class.
@@
To tackle these challenges, we propose a simple contrastive learning framework to encode discriminative features between samples for AU detection, serving as a substitution for conventional pixel-level learning. This approach offers a pathway to lightweight model development within the domain of AU detection tasks. Moreover, in response to the issue of noisy labeling in AUs, we introduce self-supervised signals to augment the noise supervision signal. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that positive examples derived from data augmentation with the self-supervised signal enhance robustness, as illustrated in Fig. 1. Additionally, to mitigate the class imbalance problem among different types of AUs, we apply re-weighting to the contrastive loss. The resultant loss, resembling the commonly utilized InfoNCE, is termed AUNCE. Consequently, modifications are made to the gradient magnitude during back-propagation in the training process. The primary contributions of our work are succinctly summarized as follows:
@@
  • We propose a novel AUNCE loss to learn contrastive feature representations for AU detection.
  • We introduce a strategy that incorporates self-supervised signal to address the issue of noisy AU labels.
@@
  • We employ an importance re-weighting strategy for minority AUs to tackle the class-imbalance problem.
@@
  • We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on two widely-used datasets, namely BP4D and DISFA.
@@ II Related Work
We will address two key aspects closely linked to our work: deep contrastive learning and contrastive learning for AU detection.
@@ II-A Deep contrastive learning
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [25], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [26], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. Despite variations in representation encoders (denoted as $f$) and similarity measures across different tasks [27, 28], the underlying concept remains consistent: training the representation encoder $f$ by optimizing a contrastive loss [29], achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
Prominent frameworks in this domain include SimCLR by Chen et al. [21], which introduces a base encoder network and a projection head trained to maximize agreement utilizing a contrastive loss. Additionally, MoCo, proposed by He et al. [30], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Chen and He [31] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Typically, these techniques are founded on the network architecture of quasi-twin neural networks. The image pairs utilized in the training process are derived from the same image, enlarged separately (positive sample pairs). The common practice involves obtaining positive samples through semantic-invariant operations on an anchor, coupled with heavy data augmentation techniques such as random cropping, flipping, image rotation, cutout, and color distortion. Conversely, image pairs are formed by enlarging different images separately (negative sample pairs).
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of numerous positive and negative sample pairs but also incorporates diverse losses for optimization. For example, Oord et al. [22] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [32]. Khosla et al. [33] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [34] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
@@
As the technology of contrastive learning advances, there is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models.
@@ II-B Contrastive learning for AU detection
While contrastive learning has garnered widespread popularity in recent years for AU detection, the predominant focus within AU detection methods has been on generative contrastive learning. Various approaches utilize a network with a parallel structure to directly compare features extracted from two or more sampled images. Wu et al. [35] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-affected facial images. Sun et al. [36] proposed an Emotion-aware Contrastive Learning framework designed to capture representations preserving sufficient AU-related information. Zhao et al. [37] devised a contrastive learning scheme aimed at creating a semantic-aware AU manifold, thereby reducing misalignment and lowering the dimensionality of facial action features. Liu et al. [38] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling. This multifaceted approach aimed to enhance the discriminative capabilities of AUs.
Moreover, certain methodologies are specifically designed to incorporate contrastive losses tailored for AU detection tasks, aiming to discern differences among AU classes or different perspectives of the same AU class. Niu et al. [39] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). In a similar vein, Suresh et al. [40] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings. Shang et al. [41] put forth a pixel-level multi-view contrastive loss, enhancing features through distinct cross-scale features. Additionally, Yang et al. [42] suggested a novel weighted multi-label contrastive learning loss tailored for AU detection.
However, AU detection methods grounded in the discriminative contrastive learning paradigm have not found widespread adoption in recent research on AU detection, particularly in the context of unlabeled video sequences. Contemporary endeavors predominantly seek to generalize similarity learning between video frames within the discriminative contrastive learning paradigm, with the objective of acquiring meaningful video-level representations [43, 44, 45, 46, 47, 48]. Nevertheless, these general self-supervised contrastive learning approaches generate representations that may not be specifically tailored to AU detection, leading to debates regarding the existence of a universal representation applicable to all downstream tasks [49]. Despite the inherent temporal coherence and evolution characteristics present in unlabeled video sequences, Li et al. [23] have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos. However, the practical application of the discriminative contrastive learning paradigm for static images to AU detection frameworks remains theoretical, and the potential substitution of traditional generative cross-entropy loss by such learning methods is a subject of ongoing discussion.
@@ III Method
In the Section. III, we mainly address three main challenges: 1. How to incentivize the encoder to capture differential information among AUs rather than pixel-level information, enabling a lightweight encoder through discriminative contrastive learning paradigm, instead of the traditional cross-entropy classification paradigm. 2. How to solve the class imbalanced problem of each AU type to improve the accuracy of our framework. 3. How to mitigate the problem of AU labels tainted by noise and errors, enhancing the robustness of our models to erroneously annotated data and thereby improving generalization performance of the models.

### v3

@@ Abstract
Facial action unit (AU) detection has long encountered the challenge of detecting subtle feature differences when AUs activate. Existing methods often rely on encoding pixel-level information of AUs, which not only encodes additional redundant information but also leads to increased model complexity and limited generalizability. Additionally, the accuracy of AU detection is negatively impacted by the class imbalance issue of each AU type, and the presence of noisy and false AU labels. In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection. To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples. Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs. This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels. Our experimental assessments, conducted on four widely-utilized benchmark datasets (BP4D, DISFA, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
@@ I Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4, 5, 6, 7, 8, 9] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction [10], emotion analysis [11, 12], and car driving monitoring [13].
In recent years, there has been a prevailing trend in supervised frameworks for AU detection [14, 15, 16, 17, 18, 19, 20]. These studies have predominantly utilized the binary cross-entropy loss to strictly align predicted and ground truth labels, which is often unnecessary for classification tasks. This overemphasis on label alignment imposes excessive demands on the fitting capacity of models, consequently increasing the risk of overfitting. Moreover, these methods commonly focus on encoding entire pixel-level information of AUs, disregarding the fact that capturing subtle feature differences when AUs activate is sufficient. As a result, these learning methods place significant demands on the model complexity to encode excessive redundant information, ultimately limiting the generalization ability of the model.
Figure 1: Overview of our proposed contrastive learning framework.
@@
Representation learning approaches rooted in contrastive learning have demonstrated substantial promise [21, 22, 23, 24, 25] for classification. This approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar [26]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [27]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [14]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [28]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
@@
To address these challenges, we propose a contrastive learning framework designed to encode discriminative features for AU detection, as depicted in Fig. 1. This approach serves as a replacement for traditional pixel-level learning methods, which also offers a pathway to the lightweight model development within the field of AU detection. The resultant loss is termed AUNCE, resembling the commonly utilized InfoNCE. Moreover, to mitigate the class imbalance issue of each AU type, we apply a negative sample re-weighting strategy, making modifications to the gradient magnitude of minority and majority class samples during back-propagation by adding importance weight to the denominator of the resultant loss. Additionally, in response to the issue of noisy and false labeling in AUs, we introduce a positive sample samplying strategy to augment the supervision signal through self-supervised signals. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that more robust positive samples are derived from data augmentation with the self-supervised signal. The primary contributions of our work are succinctly summarized as follows:
@@
  • We propose a contrastive learning framework to learn discriminative feature representations for AU detection, and the resultant loss is named AUNCE.
@@
  • We employ a negative sample re-weighting strategy to tackle the class imbalance issue of each AU type.
  • We introduce a positive sample samplying strategy to address the issue of noisy and false AU labels.
@@
  • We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on four widely-used datasets, namely BP4D, DISFA, GFT and Aff-Wild2.
@@ II Related Work
We provide a comprehensive overview of two areas closely connected to our research: deep contrastive learning and contrastive learning for AU detection.
@@ II-A Deep contrastive learning
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [30], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [31], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. Prominent frameworks in this domain include SimCLR by Chen et al. [26], which introduces a base encoder network and a projection head trained to maximize agreement utilizing a contrastive loss. Additionally, MoCo, proposed by He et al. [32], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Seyfi et al. [33] and Zhao et al. [34] have both proposed improved contrastive learning frameworks based on MoCo. Chen and He [35] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Despite variations and errors in representation encoders (denoted as $f$) and similarity measures across different tasks [36, 37], the underlying concept remains consistent: training the representation encoder $f$ by optimizing a contrastive loss [38], achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. For example, Oord et al. [27] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [39]. Khosla et al. [40] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [41] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
@@
There is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models.
@@ II-B Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task [32] for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [42, 43, 44, 45, 46]. For example, Niu et al. [42] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Liu et al. [45] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling. Additionally, Chang et al. [46] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder.
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [47, 48, 49]. for example, Wu et al. [47] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-activated facial images. Chen et al. [49] introduced a supervised hierarchical contrastive learning method (SupHCL) aimed at enhancing AU detection by fostering knowledge consistency across distinct facial images and various AUs. In a similar vein, Suresh et al. [48] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
Different from the above static image-based approaches, some methods predominantly seek to generalize similarity learning between video frames within the discriminative paradigm [50, 51, 52, 53, 54, 55]. One of the most representative methods is Li et al. [28] proposed, which have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos.
In summary, contrastive learning has demonstrated promising results in various tasks. However, its direct application to the task of detecting subtle feature differences when AUs activate presents challenges: noisy and erroneous annotations, class imbalance in AUs, and difficulty in distinguishing hard negative samples.
@@ III Method
In Section III, we present a contrastive learning framework specifically designed for the task of AU detection, aiming to incentivize the encoder to learn differential information between AUs from each sample pair, rather than focusing on pixel-level information. Subsequently, we confront two additional challenges: 1. Tackling the issue of class imbalance pertaining to each AU type in order to enhance the accuracy of our framework. 2. Addressing the problem of noisy and false labels, aiming to bolster the robustness of our models against erroneously annotated data.
III-A Contrastive learning framework
For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss. This loss function measures the disparity between predicted labels and ground truth labels. However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates. Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [14] denoted as ${L}_{wce}$
${L}_{wce}=-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}[y_{i}log\hat{y}_{i}+(1-{y}_{i})log(1-\hat{y}_{i})],$ (1)
where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability. This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones. $w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training. It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set. $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
However, the classification methods based on cross-entropy loss suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output $\hat{y}$ must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a contrastive learning framework for AU detection. As shown in Fig. 1. It comprises a positive sample mining module that integrates self-supervised signals with supervised signals, aiming to address the issue of noisy labels resulting from subtle differences in AUs. Additionally, a negative sample weighting module is included to tackle class imbalance and hard negative sample mining challenges. These modules will be discussed in detail in Sections III-B and III-C, respectively. The resultant loss is formulated as:
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\omega_{j}\cdot\text{sim}(f(x),f(x^{-}_{j}))}]\},$
where $f(\cdot)$ serves as an encoder, mapping the features to a $d$ dimensional vector, represents the feature representation of a specific AU. The positive sample of $x$, denoted as $x^{+}$, corresponds to samples belonging to the same class as $x$. Specifically, they are derived from three sources: samples from the same class, data augmentation of the anchor point, and feature augmentation of samples from the same class. $x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$. $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example. The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ and ${p}^{-}$ respectively denote the class conditional distribution of positive and negative samples of $x$. Inspired by Eq (1), we introduce $w_{i}$ in Eq (2) to mitigate the influence of loss values related to commonly activated AUs in the training set. Additionally, we incorporate $\omega_{j}$ to re-weight the significance of each negative sample.
The contrastive loss ${L}_{aunce}$ distinguishes itself from weighted cross-entropy loss ${L}_{wce}$ in the following aspects: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label $y$, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the features of samples from different classes. This incentivizes the encoder to capture the differential information among different class samples, without the need to encode pixel-level information, enabling a lightweight encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.
To ensure a fair comparison with conventional pixel-level learning paradigms, we utilize the network structure from the first training stage of MEGraph [56] as our contrastive learning encoder, as illustrated in Fig. 2. The input to the encoder consists of an contrastive larning image pair consisting of an anchor and a random sample, and the encoder is optimized by AUNCE only. Our backbone adopts Swin-transformer [57], and the AUs Relationship-aware Node Feature Learning (ANFL) module is composed of two integral components: the AU-specific Feature Generator (AFG) and the Facial Graph Generator (FGG). The AFG independently generates a unique representation for each AU. Building on these representations, the FGG constructs an optimal graph for each facial display to accurately recognize all target AUs. To accomplish this, the FGG ensures that the AFG encodes task-specific associations among AUs into their respective representations.